# Identifying Reliable Predictions in Detection Transformers

### v1

@@ 6 Experiments
To demonstrate the effectiveness of the proposed method, we conducted experiments using four DETR variants: UP-DETR [3], Deformable-DETR (D-DETR), Cal-DETR [17], and DINO. Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set. The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD). Experimental details can be found in Appendix E.
6.1 Effectiveness of OCE
To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration error across four DETR models. As baselines, we measure ECEs using the top-$100$ detections (best AP), a standard confidence threshold of 0.3, and the threshold that minimizes LRP on the validation set. In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set. For fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set. We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the Pearson correlation coefficient to the other scoring schemes across different datasets.
As shown in Table 2, our OCE metrics maintain a strong correlation with the optimal ECE metrics across different datasets. This evidence supports the reliability of our OCEs in ranking models by their overall calibration quality on positive sets—which downstream applications are more likely to utilize instead of the top 100 predictions. Notably, LA-ECE0 with a non-optimal positive set shows an inverse correlation with the rankings estimated on the optimal positive set. Meanwhile, from a model ranking perspective, empirical evidence suggests that solely adopting LRP does not align with calibration qualities. Nonetheless, the ranking by D-ECE and OCE scores measured at the optimal LRP threshold — denoted by D-ECE (LRP) and OCE (LRP), respectively — show a strong correlation to those ranked from the optimal positive set; thus, the method of using LRP to identify a positive set [11] appears acceptable. Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
Another important attribute of OCE is that it can not only be used to compare different models but is also capable of identifying the most reliable configuration of the same model. As illustrated in Figure 3, the confidence vs. OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$. This is particularly notable because it reflects the practical choices in many calibration studies to date [17, 11], affirming the reliability of using OCE.
Using a fixed separation threshold of $0.3$ is often a good approximation for distinguishing between optimal positives and negatives (See further discussion in Appendix G). However, trained or post-calibrated DETR models may have varying optimal confidence thresholds (for example, low-temperature calibrated models are likely to have higher optimal thresholds). Therefore, using a fixed threshold can pose a potential risk.
The biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed. While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference. Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
Table 3: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (6), with different models. $\mathsf{Perf}$ is evaluated using AP (Average Precision)[14]. For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), and the proposed OCE-based post-processing scheme, as indicated within the parentheses. The strongest correlations are highlighted in bold, while negative correlations appear in red. First, the proposed method achieves the highest correlation with image-level reliability. In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$).
$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.684 0.709 0.695 0.601 0.633 0.662 0.634 0.612 0.657 0.672 0.643
Ours $\mathsf{Conf}^{+}$ (OCE) 0.442 0.477 0.547 0.512 0.237 0.355 0.393 0.370 0.228 0.291 0.270 0.274
$\mathsf{Conf}^{-}$ (OCE) -0.633 -0.581 -0.571 -0.585 -0.365 -0.389 -0.379 -0.394 -0.245 -0.325 -0.253 -0.283
$\mathsf{ContrastiveConf}$ (OCE) 0.585 0.570 0.597 0.587 0.334 0.432 0.431 0.438 0.287 0.362 0.302 0.327
6.2 Quantifying Image-Level Reliability
To address the third research question (RQ3), we compare different methods for quantifying per-image reliability using different separation techniques. We measure the Pearson correlation coefficient (PCC) of each method with the $\mathsf{ImReli}$ computed based on AP metrics. Additionally, we evaluate the proposed contrasting approach against methods that utilize only positive or negative queries. Empirical results are presented in Table 3 and further in Appendix F. The key takeaways are as follows:
First, as demonstrated throughout our analysis, distinguishing between positive and negative queries is crucial for uncertainty quantification in DETR. Specifically, when confidence scores are averaged without distinguishing positive queries (i.e., $\mathsf{Conf}^{+}$ w/ top-$100$), the correlation becomes negative. This occurs because the majority of DETR’s predictions are actually negative (i.e., $|\mathcal{T}_{\bm{x}}|\ll|\hat{\mathcal{T}}_{\theta}(\bm{x})|$), and that confidence scores for negative queries are inversely correlated with $\mathsf{ImReli}$.
Second, the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability. Interestingly, the absolute value of $\mathsf{Conf}^{-}$ occasionally outperforms $\mathsf{ContrastiveConf}$ when a non-optimal separation scheme is applied; however, its performance varies across different settings. This variability underscores the robustness of our contrasting approach, which leverages the strengths of both $\mathsf{Conf}^{+}$ and $\mathsf{Conf}^{-}$.
Table 4: Comparison of different post-processing schemes on OCE (lower is better) and PCC (higher is better) between $\mathsf{ContrastiveConf}$ and image-level reliability in Cal-DETR. The thresholding approach consistently yields the best performance.
Schemes COCO City Foggy
OCE PCC OCE PCC OCE PCC
Optimal 0.425 0.695 0.463 0.634 0.513 0.643
Thr 0.454 0.594 0.448 0.464 0.482 0.325
Top-$k$ 0.502 0.527 0.453 0.411 0.492 0.298
NMS 0.555 0.574 0.582 0.373 0.598 0.239
6.3 Comparative Study on Post-Processing
Lastly, we present a comparative analysis across different post-processing schemes for identifying positive queries in DETR, from the reliability perspective. We assess OCE for each predicted positive to evaluate the impact of various separation schemes on calibration performance.
For the separation methods, we compare the following approaches: (1) applying a threshold on the confidence score, (2) selecting the top-$k$ predictions, and (3) utilizing NMS. For each method, we evaluate OCE by varying the hyperparameter (e.g., confidence threshold, $k$, and IoU threshold, respectively) and compare with the score on the optimal positive. Likewise, we assess the correlation to image-level reliability for each separation scheme. For each scheme, the best hyperparameter is chosen using the validation set. A summary of the results is provided in Table 4, with the comprehensive results available in Appendix G.
The thresholding approach achieves the smallest OCE and the best image-level UQ performance. In addition, the resulting OCE and correlation is the most comparable to the optimal score, especially for the in-distribution test dataset. The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$. The best $k$ ranges from $10$ to $20$, corresponding to the average number of objects per image in the COCO dataset. On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times. Nonetheless, a performance gap remains compared to optimal matching, especially on out-of-distribution datasets. Addressing this gap is left for future work.
7 Conclusion and Limitations
The main contribution of our work is an in-depth analysis of the importance of post-processing in DETR frameworks. Furthermore, this paper demonstrates the impact of the post-separation method on the reliability of their predictions, and proposes a novel UQ method for quantifying image-level reliability in DETR.
The primary limitation of our work is that our analysis is confined to standard separation schemes, such as thresholding, without proposing a novel separation technique. We conjecture that leveraging information from intermediate layers, rather than solely relying on final predictions, can achieve more robust separation because positive and negative queries follow different evolutionary patterns. One potential approach is to enable the model to explicitly output matching probabilities by introducing an auxiliary loss during training, as proposed by Oksuz et al. [19].
Acknowledgments
Appendix A Related Work
Uncertainty Quantification.
Shelmanov et al. [27] and Vazhentsev et al. [29] focus on uncertainty quantification in general Transformer architectures, building upon the MC dropout approach proposed by Gal and Ghahramani [7]. Specifically, they apply a modified MC dropout that prioritizes sample diversity over randomness to improve sample efficiency. However, these methods are not directly applicable to object detection problems because, unlike classification tasks, object detection involves predicting an unordered set. Consequently, it is questionable whether applying MC dropout to each output is appropriate, as the permutation of predictions could be completely shuffled after applying MC dropout.
Several studies focus on out-of-distribution (OOD) identification in object detection models. For example, Li et al. [13] propose a built-in OOD detector to isolate OOD data for human review. This approach can determine OOD scenarios involving both unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty) by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions. Du et al. [5] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data. Similarly, Oksuz et al. [19] employ an auxiliary detection model capable of expressing its confidence. Other works, including Du et al. [4] and Wilson et al. [30], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
To the best of our knowledge, the aforementioned existing uncertainty quantification techniques primarily focus on prediction-level analysis. Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection model. In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs. Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
Calibration.
Previous studies primarily addressed OOD detection to assess the prediction-level trustworthiness of DETR. They aim to introduce a framework that does not rely on the model’s confidence scores, which can be either overconfident or underconfident, especially for OOD data points. Simultaneously, additional research endeavors to determine the model’s overall reliability in its confidence scores. More specifically, these studies assess whether the model’s confidence scores align with their actual accuracy. In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy. To evaluate this alignment, they measure the Expected Calibration Error (ECE) by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin. However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes. Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction. To address these challenges, D-ECE [10] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an IoU threshold (commonly set at 0.5 or 0.75). On the other hand, LaECE [18] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
However, as demonstrated throughout our paper, DETR exhibits poor calibration on optimal negative predictions, while in practice, users prefer to utilize only predicted positive samples (e.g., high-confidence predictions). Therefore, if we rank the model’s calibration quality based on the entire prediction set, these rankings are highly biased and thus unreliable. To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$. As detailed in Appendix G, a threshold of $0.3$ approximates the optimal positive predictions in many DETR variants; thus, using such a fixed threshold may be practically acceptable. However, this threshold is not guaranteed to be $0.3$ for other models like UP-DETR, which introduces potential risks. In response, our paper addresses the importance of identifying a reliable subset of predictions and proposes the Object-level Calibration Error (OCE) metric, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
Broad Implications.
The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used, which has several important implications. For example, when building an auto-labeling system using data-driven detection models like DETR, it is crucial to consider that the model may not always provide accurate predictions. As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler. Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image. Furthermore, by providing a comprehensive analysis for identifying a reliable subset of predictions, our study offers users a systematic and trustworthy method for post-processing the outputs of DETR. We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
Figure 4: A diagram of the DETR architecture. An input image is first processed through a CNN backbone to generate a 2D feature representation. This representation is then passed to the Transformer encoder, which extracts feature vectors. These feature vectors are sent to the decoder, which receives $M$ learned object queries together. The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities. Please refer to the original papers for details.
Appendix B Bipartite Matching: Positive & Negative Queries
Since the number of queries in DETR, $M$, is much higher than the number of ground truth objects, DETR matches each ground truth object with the corresponding best model prediction during its training. To compute this optimal (i.e., ground truth) matching for the predictions in a given image, a bipartite matching algorithm is applied. More specifically, a matching cost between each pair of a given prediction and a ground truth object is defined as follows:
$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (12)
where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$. $\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [24] loss between bounding boxes. After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm.
In this paper, we then refer to those matched queries as positive predictions and the remaining unmatched queries as negative predictions. More details on the bipartite matching process and Hungarian algorithm can be found in [1].
Appendix C Further Visualizations of How DETR Generates Predictions
The predictions across all six decoder layers for each of the images presented in the main paper are provided in Figure 5. Another noticeable example is provided in Figure 6, where predictions for two different objects, a dog and a laptop, are shown for the same low-reliability image. In Figure 6(a), for the laptop object, all of the queries start off with low confidence and remain that way over the course of the layers. However, the class predictions gradually shift from dining table and person to laptop, and result in an overall low confidence score despite a correct final prediction (laptop).
However, Figure 6(b) indicates that even in a low-reliability image there can still be reliable predictions. For the dog object, the model is very confident in its predictions from the start, with many overlapping prediction sets in the same area and with matching classes. While there are two dominant predictions in the first layer, DETR focuses on a single prediction (query 0) while the confidence for query 1 is gradually reduced; this aligns with the trend observed in the reliable examples shown earlier.
Appendix D Performance Metrics
D.1 Expected Calibration Error (ECE) Metrics
Detection Expected Calibration Error (D-ECE) [10] is a calibration metric that specifically focuses on calibrating the confidence of an object detection model to align with precision. The confidence space is divided into $J$ bins, and then D-ECE is computed as follows:
[…]
where $\hat{\mathcal{D}}$ and $\hat{\mathcal{D}}_{j}$ are the set of all detections and the detections in the $j$-th bin, and $\bar{p}_{j}$ and precision$(j)$ are the average confidence and the precision of the detections in the $j$-th bin. D-ECE identifies true positive (TP) samples by confirming the accuracy of class labels and ensuring that the IoU surpasses the threshold $\tau$. Following the standard protocol [11], this work measures D-ECE using thresholds $\tau=0.5$ and $\tau=0.75$, and then averages the scores.
To account for both localization error and classification precision, LaECE [19, 11] calibrates model confidence to the product of precision and the IoU between the bounding box predictions and ground-truth boxes. LaECE is computed in a class-wise manner to combat class imbalance, as follows:
$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}|\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)|$ (14)
where $c$ refers to each class and $\bar{\text{IoU}}^{c}(j)$ refers to the average IoU of $\mathcal{\hat{D}}^{c}_{j}$. Similar to D-ECE, LaECE requires an IoU threshold, $\tau$. In contrast, LaECE0 proposes to set this threshold $\tau$ to 00 and instead set $\bar{\text{IoU}}^{c}(\cdot)=0$ for the false positive (FP) samples. This leads to a reduction of the calibration criterion of LaECE as follows:
$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}|\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)|$ (15)
D.2 Performance Metrics
$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
where $p(r)$ is calculated by the maximum precision at recall level $r$.
[…]
D.3 Impact of Separation Schemes on Metrics
The trends presented in Figure 3 also hold for other datasets and models. Figure 7 shows the performance of the model according to different metrics with various confidence thresholds (i.e. post-processing schemes). As illustrated in Section 4.1, the optimal AP is achieved when the threshold is set to $0.0$. In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence. However, as discussed earlier, using the entire or small subset carries a high risk of including uncalibrated negatives or having missing detections, leading to unreliable decisions in practical applications and diminished interpretability.
Appendix E Experiment Setup
To compute $\mathsf{ImReli}$ for each image, we use the same COCO evaluator [14] to obtain image-wise AP scores. However, rather than passing in the entire image set, the predictions for each image are passed in individually.
Appendix F Further Experimental Results
F.1 Effectiveness of OCE
Table 5 presents the correlations between the rankings provided by three different calibration metrics on the optimal positive set and various methods. The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—–are highly correlated. This indicates that all three metrics effectively capture the notion of calibration quality. Additionally, these metrics show a decent correlation with AP metrics. It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
Furthermore, LRP, D-ECE, and LA-ECE0 exhibit negative correlations with calibration quality when used alone. This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
Finally, D-ECE exhibits a high correlation when paired with either a fixed threshold or a proper adaptive threshold (i.e., using LRP and our OCE). This is because DETR’s optimal positive threshold is empirically found to be around $0.3$ (See Table 7.). Consequently, OCE aligns well with the experimentally approximated threshold and the LRP metric in identifying effective post-processing schemes. This alignment further supports the robustness and effectiveness of our method.
F.2 Quantifying Image-Level Reliability
This section provides an extended version of Table 3 with more methods such as the proposed method but with a fixed threshold. Furthermore, we show more results with different $\mathsf{Perf}$ metrics by evaluating it with AP50 and AP75 (average precision with Iou threshold 50% and 75%). The results are shown in Table 6.
As illustrated, the observation is consistent across different settings. Another notable observation is the comparison between the fixed threshold versus our adaptive threshold selection using OCE. Although the proposed fixed-threshold approach ($0.3$) occasionally outperforms the adaptive method, the performance gap remains marginal. Moreover, the fixed threshold consistently exhibits a lower correlation with UP-DETR, for which the optimal threshold has been empirically determined to be approximately $0.5$ ($0.45$ by OCE and $0.55$ by LRP). This indicates that although the OCE approach may not achieve threshold selection that best approximates the optimal positives, it ensures the robustness of uncertainty quantification performance. In contrast, fixed thresholding can degrade performance when the chosen value deviates from the optimal threshold.
Finally, we perform an ablation study on the scaling factor ($\lambda$), with the results presented in Figure 8. The study shows that the optimal scaling factor lies within the range of $5.0$ to $10.0$. Notably, this range remains effective even for out-of-distribution datasets, such as Cityscapes and Foggy Cityscapes. This consistency suggests that the scaling factor selected using the validation set (e.g., COCO) generalizes well across datasets. Furthermore, it demonstrates that $\mathsf{ContrastiveConf}$ with $\lambda>0.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$).
Appendix G Comparative Study on Post-Processing
In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality. First, we vary the hyperparameter of the standard post-processing processes. In this first analysis, to exclude the dependency of the final performance on OCE, we don’t use OCE as a selection criterion. Rather, we choose the best hyperparameter on the validation set for each setting and compare their best-possible performances. The results are shown in Table 7 .
As illustrated, we can observe that the confidence thresholding approach outperforms the top-$k$ and NMS approaches. Top-$k$ achieves the minimum OCE with top-$20$. Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable. Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions.
Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 8. We reconfirm the effectiveness of confidence thresholding; for Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level UQ applications. Moreover, the correlation is negative especially when NMS is applied along with the $\mathsf{Conf}^{+}$ framework. This is because, while most of the optimal positive queries are likely to be included with NMS, a substantial number of optimal negative queries remain within the final subset, often outnumbering the positive ones by several times. As a result, applying NMS leads to an inaccurate reliability assessment. However, many schemes, including NMS, achieve significant improvement when used with the proposed contrastive framework. This is because even if the post-processing scheme is inaccurate, the framework can robustly perform by leveraging the negative queries.
Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter. As depicted in Figure 10, the empirical results highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance. We also highlight this claim with an exemplary visualization in Figure 9.
Table 5: Comparison of the effectiveness of the proposed OCE and baseline metrics, focusing on their ranking correlation with the rankings given by the models’ calibration quality on optimal positive predictions (denoted by the superscript *). For each metric, various post-processing schemes (i.e., ways to identify positive predictions) are applied and indicated within parentheses. Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red. Our OCE-based rankings consistently provide strong correlations with optimal ranking and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.

### v2

@@ IV-C Numerical Analysis
Setup. To provide quantitative support for our claim regarding the varying levels of reliability across predictions, we conducted experiments using four DETR variants: UP-DETR [35], Deformable-DETR (D-DETR), Cal-DETR [12], and DINO. Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set. The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD). Further details can be found in the Appendix.
Object-level Calibration Error. We evaluate and compare the object-level calibration errors on the optimal positive predictions and the optimal negative predictions. The optimal positives and negatives are determined using the previously mentioned bipartite matching algorithm.
As illustrated in Figure 3, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy. Furthermore, as shown in Figure 4a, the optimal positive predictions are significantly better calibrated than the negative ones. This trend is evident across different state-of-the-art DETR variants and datasets, all of which exhibit poor calibration quality for optimal negative predictions.
Correlation to Image-Level Reliability. We compute $\mathsf{ImReli}$ for each image by using the same COCO evaluator [18] to obtain the image-wise AP score for $\mathsf{Perf}$ in Equation 1, passing the predictions and annotations for each image individually rather than the entire image set. Then, we measure the Pearson correlation coefficient (PCC) between the $\mathsf{ImReli}$ and the average confidence score of the optimal positives and the optimal negatives, respectively.
As shown in Figure 4b, the average confidence scores of the optimal positive predictions exhibit a moderately strong positive correlation with image-level reliability (i.e., the model’s actual average precision on that image) across different models. Notably, the average confidence scores for negative predictions are inversely correlated with image-level reliability. This finding reinforces our claim, highlighting the importance of distinguishing between positive and negative predictions, especially when conducting image-level UQ in DETR. Further visualizations for the other datasets are provided in the Appendix.
V A Systematic Framework for Identifying Positive Predictions
As illustrated so far, using positive predictions is crucial for the reliable use of DETR. Nonetheless, due to the lack of annotations, it is not feasible to identify either the ground truth or the optimal positive predictions during the inference stage. Therefore, an alternative systematic framework is essential not only for improving its interpretability but also for ensuring reliability. Given that the optimal matching achieves a small calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
Figure 5: Impact of confidence threshold selection on various performance and error metrics in Cal-DETR evaluated on COCO. Higher scores are preferable ($\uparrow$), while lower scores are preferable ($\downarrow$).
V-A Limitations of Existing Metrics
To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and localization recall precision (LRP) [21, 36], for identifying positive predictions. Specifically, we sweep the threshold for the confidence score to generate different subsets of DETR predictions (i.e., the predicted positives). We then assess the performance of each metric for these subsets and determine the threshold value that yields the highest performance. Analytical results using Cal-DETR are illustrated in Figure 5. For definitions of AP and LRP and additional results on the other DETR models, please refer to the Appendix.
Primarily, as noted in several studies [21, 19, 20, 36] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$ because AP does not penalize harshly for having low confidence predictions. However, as discussed earlier, using the entire prediction set carries a high risk of including uncalibrated negatives, leading to unreliable decisions in practical applications. Furthermore, using hundreds of predictions diminishes the interpretability of the model.
In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence. This is a structural pitfall that prediction-level ECEs commonly face [20]. Since ECEs do not penalize missed detections (i.e., false negatives), they could achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions. Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when paired with a large confidence threshold.
On the other hand, LRP, a localization-focused performance metric, could be used instead. However, LRP is not designed as a calibration metric so does not explicitly consider calibration error. As a result, the model ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities, as shown in the following numerical analysis.
V-B Proposed Metric: Object-Level Calibration Error
Notation. For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider a set of ground truth object annotations $\mathcal{D}_{i}=\{d_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}}$, and a set of DETR predictions: $\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})=\{\hat{d}_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries. Here, $l_{i,j}\in[1,C]$ and $\hat{\bm{p}}_{i,q}\in\{0,1\}^{C}$ represents the ground-truth label and predicted probability distribution over $C$ classes, respectively, while $\bm{b}_{i,j}$ and $\hat{\bm{b}}_{i,q}\in\{0,1\}^{4}$ corresponds to the ground truth and predicted scaled bounding box, respectively.
Definition 2.
Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{A}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{A}\circ\hat{\mathcal{D}}_{\theta}$. We define an object-level calibration error (OCE) as the average Brier score per object:
$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}$ (6)
$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\bar{p}_{i,j}[c]\big{)}^{2}$ (7)
$\bar{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (8)
where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $d_{i,j}$ and we propose two variants:
$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\delta\}~{}$ $(\mathsf{OCE_{ENS,\delta}})$ (9)
$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmax}_{q}\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (10)
The difference is that $\mathsf{OCE_{ENS,\delta}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching. When $|\mathcal{Q}_{i,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$. Following [18, 20], we use IoU thresholds of $\delta=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
The introduced calibration error has two desirable characteristics. First, the prediction set achieves the lowest calibration error when the predictions are well-calibrated to the respective closest ground truth objects. Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects. This is achievable due to the primary difference between our OCE and ECEs: while ECEs are evaluated based on selected predictions, OCE is assessed along the ground truth objects. This ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics. Thus, this metric can effectively assess not only whether the given subsets $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compared to existing metrics.
V-C Numerical Analysis: Effectiveness of OCE
Table 1: Comparison of the effectiveness of the proposed OCE and exisiting metrics, focusing on their correlation with the models’ overall calibration quality on optimal positive predictions (denoted by the superscript *). For each metric, different post-processing schemes (i.e., methods for identifying positive predictions) are applied and specified in parentheses. These include using the top 100 predictions or choosing predictions with confidence scores above the threshold that optimizes the respective metric. Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red. Our OCE consistently provides strong correlations with optimal calibration errors and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.